OpenStudy (brittany):

Use the rules for polynomial operations to simplify the following expression. Show all your computations and explain the steps you take to arrive at your answer. a. (3x-2)(6x+1) b. 12t(t + m)(t - 3m)

jimthompson5910 (jim_thompson5910):

a) \[(3x-2)(6x+1)\] \[(3x)(6x)+(3x)(1)+(-2)(6x)+(-2)(1)\] \[18x^2+3x-12x-2\] \[18x^2-9x-2\] So \[(3x-2)(6x+1)=18x^2-9x-2\] ---------------------------------------------------- b) \[12t(t + m)(t - 3m)\] \[12t\left((t)(t)+(t)(-3m)+(m)(t)+(m)(-3m)\right)\] \[12t(t^2-3mt+mt-3m^2)\] \[12t(t^2)+12t(-2mt)+12t(-3m^2)\] \[12t^3-24mt^2-36m^2t\] So \[12t(t + m)(t - 3m)=12t^3-24mt^2-36m^2t\]